# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

### v2

@@ V-C Ablation Studies
Effect of the Tri-plane Selection. We first compare our method with the single-plane methods to evaluate the effectiveness of the tri-plane. For the single-plane method, we choose one plane as the main plane and only track along the chosen plane. The results on the BvEM-Macaque volume are shown in Tab. III. We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane. This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others. Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other. Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.
TABLE III: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
@@
Figure 9: Histogram of the tri-plane selection.
Figure 10: Dynamics of the tri-plane selection.
@@
TABLE IV: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
@@
Effect of the Recursive Redirection Strategy. To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive seed sampling by considering potential turning points. To validate its effectiveness, we report the results on the BvEM-Macaque volume in Tab. IV where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included. Compared to the baseline without recursive seed sampling “naive” and dense seed sampling “dense”, the proposed method achieves the best performance with less than 10% running time increase. One naive baseline is to remove the recursive seed sampling component and not consider any potential turning points. This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance. Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes. Therefore, the performance is limited. The proposed method achieves the best performance with less than 10% running time increase on the BvEM-Macaque dataset. Unfortunately, it significantly increases the computation cost. Compared to our method, DenseSAM’s running time is 5.97 times longer since it needs to segment 3 planes for every tracking step. However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse. This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
TABLE V: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
@@
Quality Analysis. We delve deeper into the plane dynamics in Fig. 11 and 11. The majority of the selected plane predominantly tracks along the y-axis. This observation aligns with the experimental results presented in Tab.III, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis. This is because the blood vessel flows mainly along the y-axis in the test volume. In Fig. 11, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses. The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
Other Ablations. To further improve the efficiency, we use MobileSAM instead of the original SAM in practice. As shown in Table V, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed. Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.Fig. 8 illustrates the performance with varying hyperparameters $\eta$ and $\tau$.

### v3

@@ 5.3 Ablation Studies
Effect of the Tri-plane Selection. We first compare our method with the single-plane methods to evaluate the effectiveness of the tri-plane. For the single-plane method, we choose one plane as the main plane and only track along the chosen plane. The results on the BvEM-Macaque volume are shown in Tab. 3. We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane. This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others. Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other. Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.
Table 3: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
@@
Table 4: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
@@
Effect of the Recursive Redirection Strategy. To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive seed sampling by considering potential turning points. To validate its effectiveness, we report the results on the BvEM-Macaque volume in Tab. 4 where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included. Compared to the baseline without recursive seed sampling “naive” and dense seed sampling “dense”, the proposed method achieves the best performance with less than 10% running time increase. One naive baseline is to remove the recursive seed sampling component and not consider any potential turning points. This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance. Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes. Therefore, the performance is limited. The proposed method achieves the best performance with less than 10% running time increase on the BvEM-Macaque dataset. Unfortunately, it significantly increases the computation cost. Compared to our method, DenseSAM’s running time is 5.97 times longer since it needs to segment 3 planes for every tracking step. However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse. This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
Table 5: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
@@
Quality Analysis. We delve deeper into the plane dynamics in Fig. 11 and 11. The majority of the selected plane predominantly tracks along the y-axis. This observation aligns with the experimental results presented in Tab.3, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis. This is because the blood vessel flows mainly along the y-axis in the test volume. In Fig. 11, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses. The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
Other Ablations. To further improve the efficiency, we use MobileSAM instead of the original SAM in practice. As shown in Table 5, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed. Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.Fig. 8 illustrates the performance with varying hyperparameters $\eta$ and $\tau$.